OpenStudy (anonymous):

(2x)^2 (1/2y^2)^2 simplify

OpenStudy (anonymous):

(2x)^2 (1/2y^2)^2 4x^2(1/4y^4) 4x^2/4y^4= x^2/y^2

OpenStudy (anonymous):

x^2/y^4

jhonyy9 (jhonyy9):

4x2/4y4 =x2/y4

OpenStudy (anonymous):

\[(2x)^2 (1/2y^2)^2\] (2x)^2 = 2x * 2x = 4x^2 (1/2y^2)^2 = 1/2y^2 * 1/2y^2 = 1/4y^4 4x^2 * 1/4y^4 = 4x^2/4y^4 = x^2/y^4 x^2/y^4